# Towards Real-time Adaptive Anisotropic Image-to-mesh Conversion for Vascular Flow Simulations

### v1

@@ 5.2 Quantitative and Qualitative Analysis
where v is a vertex of element k and $M_{mean}$ is the interpolated metric tensor evaluated at the centroid of element k. Given that optimal edges should be unit-length, edges with length above or below one are considered to be sub-optimal. The measure for mean ratio is bounded between zero and one since it is normalized by the volume of an equilateral element. One is the optimal quality for an element’s mean ratio shape. LABEL:fig:arteries_CDT3D_MR_EL shows that CDT3D maintains good quality when generating an anisotropic mesh for the first case. Although there are 3 elements with a mean ratio measure less than 0.1 (the minimum shape measure is 0.016), the majority of elements have a high mean ratio and edge lengths close to 1. With regards to the second case, the anisotropic mesh generated by CDT3D exhibits poor quality when the surface of the input mesh is not adapted (containing about 1K elements with a mean ratio less than 0.1). However, when CDT3D is permitted to adapt the surface, the final mesh quality is much better (about 20 elements have a mean ratio less than 0.1, with the minimum shape measure as 0.042). When adapting the surface, the same parameters as those in table 1 are utilized with surface adaptation turned on. It should be noted that CDT3D converges and completes anisotropic adaptation in less time when surface adaptation is permitted (generating 90K tetrahedra in about 22 seconds when utilizing 20 cores) as opposed to its performance when restricting the method to preserve the input surface. The quality of the CDT3D mesh generated for the first case shows that CBC3D indeed generates a mesh of good quality when converting the first case image, given that surface adaptation is turned off for CDT3D and it still generates a final volume of good quality. In order to provide a volume that conforms to the boundary layer grid generated by AFLR, surface adaptation is turned off and CDT3D is therefore limited by the quality of the input surface mesh for the second aneurysm case. It should be noted that surface adaptation was attempted first, where the CDT3D-adapted surface was provided as input to AFLR. However, AFLR failed to process the CDT3D-adapted surface and generate a boundary layer grid. We also attempted surface refinement and volume generation using AFLR, to then subsequently use as a background mesh for CDT3D’s anisotropic adaptation (which could potentially be merged with a boundary layer grid generated over the AFLR-refined surface). AFLR however fails to refine the input PLC surface of the second case. Nevertheless, the majority of elements in both anisotropic meshes generated by CDT3D for the second aneurysm case (with and without surface adaptation) have a high mean ratio and edge lengths close to 1.
@@
Table 5: Smallest and largest dihedral angles in isotropic meshes generated by TetGen and CDT3D for both aneurysm cases.
@@
TetGen (iso-sizing) Smallest 1.67 -
Largest 176.73 -
TetGen (iso) Smallest 3.99 2.71
Largest 171.66 171.3
@@
6 Discussion and Future Work.
@@
Although it doesn’t generate anisotropic meshes, TetGen exhibits good quality while generating much fewer elements compared to CDT3D (for the small test cases of our evaluation). Additionally, TetGen meets the real-time requirement for small-size problems. CDT3D generates meshes with better quality but at the cost of generating more elements. Depending on the accuracy needed for CFD simulations, this tradeoff may be acceptable, particularly when combined with CDT3D’s anisotropy (seen in LABEL:fig:Mesh_Cuts_artery1 and LABEL:fig:Mesh_Cuts_artery2). Both the Build_Sizing/Metric and CDT3D methods show good scalability on a single multicore node, particularly when generating an anisotropic mesh. Due to the fact that CDT3D is designed to handle much larger problems (such as that in fig. 2 and LABEL:fig:artery2_images), its performance suffers from the overhead of parallelizing adaptation for such small meshes. Nevertheless, it also meets the real-time requirement when utilizing a full multicore node (for anisotropic mesh adaptation) and when utilizing a single CPU core for isotropic mesh generation. CDT3D exhibits much faster performance when generating isotropic meshes from the PLCs of the first and second cases.
The parallelism offered by CDT3D would be well suited for processing a geometry generated from an image of a large vascular structure (such as in LABEL:fig:artery2_images). This would likely require many more elements compared to the small test cases of this study. Consequently, additional memory would also be required, potentially necessitating the need for a distributed memory approach (such as [35]) if a sequential or multicore cc-NUMA-based method cannot generate and store such a large mesh within the memory of a single machine (millions of elements, or potentially hundreds of millions of elements which in the future may be needed for aneurysm cases requiring even finer mesh resolution, e.g., microvascular aneurysms). Table 3 shows the elements generated per second up to 20 cores by CDT3D for the first aneurysm case. As mentioned previously, the adaptation rate becomes 27K when utilizing 40 cores. For comparison, TetGen generates 20K isotropic elements per second (when using the sizing function for adaptivity). Consider an example where a vascular structure requires a mesh of 10 million elements in order to obtain the needed accuracy in a CFD simulation. Based on the above rates, TetGen would take approximately 8 minutes (or 83 minutes in the case of a 100 million element mesh) to generate an adaptive isotropic mesh while CDT3D would take approximately 6 minutes (or 61 minutes for 100 million elements) to generate an adaptive anisotropic mesh using 40 cores. Furthermore, CDT3D utilizes a similar speculative execution model (described in section 3.3) that was utilized within the I2M method in [31]. With an additional hierarchical work-stealing optimization for load balancing, this model for parallelization was tested and shown to maintain a speedup of approximately 123 when executed on a distributed shared memory architecture using 144 cores (seen in table 4 of [31]). If CDT3D were to receive a similar optimization and be executed on a similar architecture, it could potentially achieve an adaptation rate of 98K elements per second (based on the rates seen thus far in table 3), generating approximately 10 million anisotropic elements in 1.7 minutes (100 million elements in 17 minutes). CDT3D’s adaptation rate, coupled with its scalable performance and better quality, may prove beneficial when anisotropy is needed to provide better simulation results compared to isotropy (such as in [66, 16]). Future work is required to optimize and test CDT3D on a distributed shared memory architecture in order to gauge its potential and verify this performance estimation.

### v3

@@ 5.2 Quantitative and Qualitative Analysis of I2M Pipeline
Figure 12(a) compares the quality of the isotropic meshes generated by each method for case 1 and Figure 12(b) compares those generated for case 2. Table 5 shows the smallest and largest dihedral angles found in the isotropic meshes generated by each method. The CDT3D meshes contain better quality compared to the TetGen and PODM meshes. This is particularly seen in (b) and (c) of Figure 12(a) and Figure 12(b), which show that the TetGen and PODM meshes contain more elements with small and large angles. Although the AFLR boundary layer grids contain many elements with small angles, the grids have a good overall distribution of angle quality. Recall that qualitative results are examined with respect to their metric conformity for the adapted anisotropic meshes. Figure 16(a) shows that CDT3D maintains good quality when generating an anisotropic mesh for the first case. Although there are few elements with a mean ratio measure less than 0.1, the majority of elements have a high mean ratio and edge lengths close to 1. With regards to the second case, the anisotropic mesh generated by CDT3D exhibits poor quality when the surface of the input mesh is not adapted (containing more elements with a mean ratio less than 0.1). However, when CDT3D is permitted to adapt the surface, the final mesh quality is much better. When adapting the surface, the same parameters as those in Table 1 are utilized with surface adaptation turned on. It should be noted that CDT3D converges and completes anisotropic adaptation in less time when surface adaptation is permitted (generating 90K tetrahedra in about 22 seconds when utilizing 20 cores) as opposed to its performance when restricting the method to preserve the input surface. The quality of the CDT3D mesh generated for the first case shows that CBC3D indeed generates a mesh of good quality when converting the first case image, given that surface adaptation is turned off for CDT3D and it still generates a final volume of good quality. In order to provide a volume that conforms to the boundary layer grid generated by AFLR, surface adaptation is turned off and CDT3D is therefore limited by the quality of the input surface mesh for the second aneurysm case. It should be noted that surface adaptation was attempted first, where the CDT3D-adapted surface was provided as input to AFLR. However, AFLR failed to process the CDT3D-adapted surface and generate a boundary layer grid. We also attempted surface refinement and volume generation using AFLR, to then subsequently use as a background mesh for CDT3D’s anisotropic adaptation (which could potentially be merged with a boundary layer grid generated over the AFLR-refined surface). AFLR however fails to refine the input PLC surface of the second case. Nevertheless, the majority of elements in both anisotropic meshes generated by CDT3D for the second aneurysm case (with and without surface adaptation) have a high mean ratio and edge lengths close to 1.
(a) Quality statistics are shown comparing the dihedral angles of the isotropic meshes generated by TetGen, CDT3D, PODM, and AFLR (boundary layer) for the first (carotid cavernous) aneurysm case. (a) shows the full distribution of element dihedral angles. (b) shows the distribution of elements with dihedral angles between 0 and 40 degrees. (c) shows the distribution of elements with dihedral angles between 140 and 180 degrees.
(b) Quality statistics are shown comparing the dihedral angles of the isotropic meshes generated by TetGen, CDT3D, and AFLR (boundary layer) for the second (middle cerebral artery bifurcation) aneurysm case. (a) shows the full distribution of element dihedral angles. (b) shows the distribution of elements with dihedral angles between 0 and 40 degrees. (c) shows the distribution of elements with dihedral angles between 140 and 180 degrees.
@@
Table 5: Smallest and largest dihedral angles in isotropic meshes generated by each method for both aneurysm cases.
@@
TetGen (no adaptivity) Smallest 1.67 2.71
Largest 176.73 171.3
TetGen (w/ adaptivity) Smallest 3.99 -
Largest 171.66 -
AFLR (bl) Smallest 1.48 2.8
Largest 176.69 171.83
PODM (no vol. adaptivity) Smallest 4.71 -
Largest 170.15 -
PODM (w/ vol. adaptivity) Smallest 4.65 -
Largest 170.14 -
@@
(a) Shown are distributions of quality statistics for the mean ratio, (a) and (b), and edge lengths, (c) and (d), of elements within the anisotropic meshes generated by CDT3D for both aneurysm cases (in linear and logarithmic scales).
5.3 Hierarchical Load Balancing
We evaluate the effectiveness of the hierarchical load balancing technique using much larger test cases (i.e., more work per thread) than those seen in the previous section. Figure 16 compares the speedup achieved with and without the updated load balancing model on three supercomputers - Purude’s Anvil, ODU’s Wahab, and ODU’s Turing supercomputers. The purpose is to gauge any potential improvements when utilizing smaller machines like the Turing supercomputer (with older specifications) versus utilizing larger machines with newer specifications (as described in section 5.1). This strong scaling experiment is tested utilizing the CDT3D-generated isotropic volume mesh of the second aneurysm case from the previous section as input. The velocity-based metric tensor field is scaled from a complexity of about 20 thousand to a complexity of 50 million (generating approximately 100 million elements). We first observe that the hierarchical load balancing model gives a noticeable improvement on the Anvil supercomputer (up to about 5% additional speedup when utilizing 64 cores). On the other hand, the new model does not improve performance much on Turing and when utilizing 40 cores on Wahab. CDT3D’s original load balancing model performs slightly better than the new model on these smaller machines. A detailed analysis of the hierarchical load balancing model and its performance results is provided in section 6.
Figure 16: A strong scaling speedup performance comparison is given of the original (Orig) and hierarchical load balancing (HLB) versions of CDT3D when generating approximately 100 million elements.
5.4 Optimized Local Reconnection
We gauge the performance improvement given by the optimized local reconnection algorithm in CDT3D, discussed in section 4.3, by testing the updated method with both aneurysm cases on the Wahab supercomputer. We scale the complexity of the metric tensor field for the carotid cavernous aneurysm from about 180 thousand to 10 million (generating approximately 50 million elements using the CBC3D-generated volume mesh as input). We again scale the complexity of the middle cerebral artery bifurcation aneurysm case to 50 million (generating approximately 100 million elements). Threads are pinned to CPU cores for each of these executions (using CDT3D’s respective input parameter, shown in table 1). Table 6 shows the runtimes for the first (carotid cavernous) aneurysm case while table 7 shows the results for the second case. Recall that there are two phases of CDT3D’s operation pipeline, within which the local reconnection operation is executed (during mesh adaptation, MA, and during final quality improvement, QI). For both aneurysm cases, the MA local reconnection time remains about the same. This is because the mesh adaptation phase (while it does improve quality) focuses on satisfying point spacing criteria in the metric field. This means that this operation is executed over elements until little to no new points are created/inserted into the grid. On the other hand, we see a significant performance improvement for the Quality Improvement phase’s local reconnection operation (by about 3 times). When executed sequentially on Wahab, this saves 1 (for the first case) to 3 hours (for the second case) in end-to-end time compared to the original CDT3D. When utilizing 40 cores, there is still a 3x reduction in QI local reconnection time but there isn’t as big of an improvement in end-to-end time. This is because the QI local reconnection time occupies only about 15% of end-to-end runtime in the original CDT3D. This is reduced to about 5% with its optimization. The results seen specifically for the QI local reconnection time shows the performance enhancement that such an algorithm can provide for local reconnection-based methods. Figures 20(a) and 20(b) show that the quality of the meshes generated by the local reconnection-optimized CDT3D is preserved, as they maintain comparable quality to that generated by the original CDT3D.
Table 6: A runtime comparison (approximately in minutes) of the original CDT3D and its optimized Local Reconnection (LR) version is shown when generating a mesh for the first (carotid cavernous) aneurysm at 10 million complexity on the Wahab supercomputer. ‘MA’ means the Mesh Adaptation phase of CDT3D while ‘QI’ means the Quality Improvement phase of CDT3D.
Method CPU Cores
1 5 10 20 40
Original CDT3D MA Local Reconnection 232 41 21 11 7
QI Local Reconnection 111 20 10 5 2.5
End-to-end 501 94 49 25 15
Optimized CDT3D MA Local Reconnection 228 40 21 11 7
QI Local Reconnection 38 6 3 1.6 0.8
End-to-end 421 79 42 22 13
Table 7: A runtime comparison (approximately in minutes) of the original CDT3D and its optimized Local Reconnection (LR) version is shown when generating a mesh for the second (middle cerebral artery bifurcation) aneurysm at 50 million complexity on the Wahab supercomputer. ‘MA’ means the Mesh Adaptation phase of CDT3D while ‘QI’ means the Quality Improvement phase of CDT3D.
Method CPU Cores
1 5 10 20 40
Original CDT3D MA Local Reconnection 524 96 52 36 18
QI Local Reconnection 249 42 21 10 5
End-to-end 1097 210 110 67 33
Optimized CDT3D MA Local Reconnection 507 95 51 33 18
QI Local Reconnection 91 13 6 3 1.6
End-to-end 916 181 95 57 30
(a) Shown are distributions of quality statistics for the mean ratio, (a) and (b), and edge lengths, (c) and (d), of elements within the meshes generated by the local reconnection-optimized CDT3D for the first (carotid cavernous) aneurysm case at 10 million complexity (in linear and logarithmic scales).
(b) Shown are distributions of quality statistics for the mean ratio, (a) and (b), and edge lengths, (c) and (d), of elements within the meshes generated by the local reconnection-optimized CDT3D for the second (middle cerebral artery bifurcation) aneurysm case at 50 million complexity (in linear and logarithmic scales).
Finally, we show the runtimes when generating large meshes on the Anvil supercomputer by the optimized CDT3D (including both its hierarchical load balancing model and local reconnection optimization), PODM’s Function Approximation (where its $\delta$ parameter is adjusted to 0.127 for volume adaptivity and 0.25 when executed without volume adaptivity to generate more elements for the first aneurysm case), and refine [ParkRefine]. Although it is originally designed and used heavily for aerospace cases [TsolakisEvaluation], refine is ultimately a state-of-the-art open-source, distributed memory adaptive anisotropic mesh generation method. We gauge its capability to be utilized for a medical case. refine utilizes a combination of edge split, edge collapse, and point smoothing operations to generate a unit grid that adheres to a given metric field. In our evaluation, the default parameters for refine are utilized with the exception of the domain decomposition method. We use ParMETIS, as this resulted in the best runtimes for refine among its domain decomposition options. Both CDT3D and refine utilize the scaled metric tensor field(s) to generate large meshes. Threads are pinned to CPU cores for both CDT3D’s and PODM’s executions (using their respective input parameters). It should be noted that PODM’s hierarchical load balancing model is not utilized because its source code was designed and written specifically for the Pittsburgh Supercomputer Center’s Blacklight architecture. We instead utilize PODM’s random work-stealing algorithm for its load balancing [foteinos2014high]. Table 8 shows the runtimes for PODM and CDT3D when generating approximately 50 million elements for the first (carotid cavernous) aneurysm case. Table 9 shows the approximate mesh sizes generated by each method for the first case. PODM generates its isotropic meshes in less than a minute when using 96 cores while CDT3D generates its anisotropic meshes in about 5 minutes. PODM exhibits excellent end-user productivity by generating an adaptive mesh at a rate of about 1.17M elements/sec. The optimized CDT3D exhibits an adaptation rate of about 143K elements/sec for the first aneurysm case. Figure 23(a) shows the dihedral angle distribution of the large meshes generated by PODM. The method generates elements with the same relative distribution of element quality as with the smaller cases. Table 10 shows that the minimum and maximum angles in both meshes remain about the same. PODM maintains good quality when performing volume adaptivity to generate large meshes as well as the small ones.
Table 8: The time spent (approximately in minutes) when generating large-size meshes (about 50 million elements) for the first (carotid cavernous) aneurysm case on the Anvil supercomputer is shown.
CPU Cores
Methods 1 2 4 8 16 32 64 96
PODM (no adaptivity) 9.6 4.8 2.4 1.3 0.7 0.4 0.3 0.2
PODM (w/ adaptivity) 50 25 13 6.6 3.4 1.8 1.1 0.7
CDT3D (Original) 294 142 75 39 21 12 8 5.9
CDT3D (Optimized) 259 122 65 35 19 10 7.8 5.5
Table 9: Number of tetrahedra and points generated by each method when generating large meshes for the first (carotid cavernous) aneurysm case at 10 million complexity. ‘Tets’ is short for tetrahedra. ‘M’ means million.
Methods Tets Points
PODM (no adaptivity) 48.2M 8.6M
PODM (w/ adaptivity) 49.3M 12.2M
CDT3D (Original) 47.4M 8.1M
CDT3D (Optimized) 47.3M 8.1M
(a) Quality statistics are shown comparing the dihedral angles of the large (approx. 50 million element) isotropic meshes generated by PODM for the first (carotid cavernous) aneurysm case. (a) shows the full distribution of element dihedral angles. (b) shows the distribution of elements with dihedral angles between 0 and 40 degrees. (c) shows the distribution of elements with dihedral angles between 140 and 180 degrees.
Table 10: Smallest and largest dihedral angles in the large isotropic meshes generated by PODM for the first (carotid cavernous) aneurysm case
Methods Smallest Largest
PODM (no vol. adaptivity) 4.35 170.29
PODM (w/ vol. adaptivity) 4.25 170.26
Table 11 shows the runtimes for CDT3D and refine when generating approximately 100 million elements for the second (middle cerebral artery bifurcation) aneurysm case. CDT3D exhibits good scalability for both aneurysm cases. Some of the runtimes are not reported for refine because the distributed method crashes due to running out of memory when executed with 16 cores or less. Table 12 shows the approximate mesh sizes generated by each method for the second case. Figure 27(a) shows the quality of the meshes generated by both CDT3D (optimized) and refine for the second case. While the CDT3D mesh exhibits good overall quality, the mesh generated by refine exhibits better quality. However, this is at the cost of generating about 60 million additional elements compared to CDT3D to satisfy the same level of metric conformity. refine also requires much more time, as it is about 9 times slower than CDT3D when using 96 cores. CDT3D also exhibits a faster adaptation rate of about 120K elements/sec when using 96 cores while refine has an adaptation rate of about 23K elements/sec. Consequently, if the user is limited to utilizing a smaller machine (i.e., less CPU cores), refine may be suitable for predictive simulations if time is not a primary concern and high mesh quality is desired for better solution accuracy with a solver. There is a tradeoff between quality and mesh adaptation speed when compared to CDT3D.
Table 11: The time spent (approximately in minutes) when generating large-size meshes for the second (middle cerebral artery bifurcation) aneurysm case at 50 million complexity on the Anvil supercomputer is shown.
CPU Cores
Methods 1 2 4 8 16 32 64 96
CDT3D (Original) 658 325 178 93 51 28 18 13
CDT3D (Optimized) 554 265 146 81 43 24 16 12
refine - - - - - 399 131 113
Table 12: Number of tetrahedra and points generated by each method when generating large meshes for the second (middle cerebral artery bifurcation) aneurysm case at 50 million complexity. ‘Tets’ is short for tetrahedra. ‘M’ means million.
Methods Tets Points
CDT3D (Original) 98.4M 16.6M
CDT3D (Optimized) 97.9M 16.5M
refine 158.4M 27.1M
(a) Shown are distributions of quality statistics for the mean ratio, (a) and (b), and edge lengths, (c) and (d), of elements within the meshes generated by the optimized CDT3D and refine for the second (middle cerebral artery bifurcation) aneurysm case at 50 million complexity (in linear and logarithmic scales).
6 Discussion
@@
Although they only generate isotropic meshes, TetGen and PODM exhibit good quality while generating fewer elements compared to CDT3D’s anisotropic mesh generation (for the small test cases of our evaluation). Additionally, TetGen and PODM meet the real-time requirement for small-size problems (and large-size problems with PODM). CDT3D generates isotropic meshes with better quality but at the cost of generating more elements. This further showcases that local reconnection-based techniques may still offer better output mesh quality than Delaunay-based techniques. Depending on the accuracy needed for numerical simulations, this tradeoff between element count and quality may be acceptable, particularly when combined with CDT3D’s anisotropy (seen in Figures 9(b) and 9(c)). Both the Build_Sizing/Metric and CDT3D methods show good scalability on a single multicore node, particularly when generating an anisotropic mesh. Due to the fact that CDT3D is designed to handle much larger problems (such as that in Figure 5(a) or the scaled complexity benchmarks), its performance suffers from the overhead of parallelizing adaptation for such small meshes. Nevertheless, it also provides near real-time performance when utilizing a full multicore node (for anisotropic mesh adaptation) and when utilizing a single CPU core for isotropic mesh generation. CDT3D exhibits much faster performance when generating isotropic meshes from the PLCs of the first and second cases.
Several factors can be attributed to the observations made in section 5.3 about the hierarchical load balancing model and its performance. The latency introduced by remote memory accesses between numa nodes can still become a performance bottleneck on the Anvil supercomputer. This is because the machine is much larger (8 numa nodes per shared memory node of 128 cores) than Turing (2 numa nodes per shared memory node of 32 cores) and Wahab (4 numa nodes per shared memory node of 40 cores). When executing the command numactl -H at the command line on a cc-NUMA high-performance computing machine, one can see the mappings of CPU cores to numa nodes and more specifically, the distance between numa nodes. Indeed, the distance between numa nodes on Anvil is greater for some nodes than the farthest distance between numa nodes on Wahab and Turing (given that they have fewer nodes and cores than Anvil). Consequently, the hierarchical load balancing model makes an impact on a larger machine like Anvil.
To verify that the model indeed serves its purpose of reducing remote memory accesses between numa nodes, we also tested our method using Intel’s vtune profiler [^3] https://www.intel.com/content/www/us/en/developer/tools/oneapi/vtune-profiler.html. Given that it requires special access to system files, we only tested our method with vtune on the Wahab supercomputer (using 40 cores) when adapting the same artery case at 50 million complexity. Table 13 shows several statistics given by vtune’s memory access analysis of both versions of CDT3D. The local and remote access counts refer to the total local and remote accesses among the two sockets (recall that there are two sockets in a single shared memory node in these machines). The hierarchical load balancing model indeed reduces the number of remote memory accesses by almost 50%.
Table 13: The approximate memory accesses of the original CDT3D’s load balancing model vs. the hierarchical load balancing (HLB) model is shown - ’M’ means million.
Original HLB
Local Memory Accesses 11.4M 12.6M
Remote Memory Accesses 50.7M 26.4M
Remote Cache Accesses 46.4M 31.5M
There is however an overhead from determining an idle thread with the shortest numa node distance every time a busy thread attempts to give work to another. The worst case time complexity for each thread (based on algorithm 1) is $O(t\log(b))$ where t is the number of idle threads and b is the number of buckets assigned to this particular thread. When utilizing 40 cores on Wahab, this accounts for approximately 6% overhead. Coupled with the fact that there are few numa nodes on this machine (and consequently, a less expensive distance between its 4 numa nodes compared to Anvil’s 8 numa nodes), the hierarchical model barely offers improvement in terms of speedup on Wahab. While offering an improvement on Anvil, it should be noted that the hierarchical load balancing model does not offer as much an improvement as seen in its implementation from [foteinos2014high]. Simply put, this is attributed to the difference in the newer specifications of Anvil as opposed to the Pittsburgh Supercomputing Center’s Blacklight machine utilized in [foteinos2014high]. Some of these differences include (1) DDR4 memory supported by Anvil (6 memory channels per CPU, offering higher memory bandwidth and lower latency) as opposed to the Blacklight’s DDR3 memory (4 memory channels per CPU) and (2) the Ultra Path Interconnect (UPI) links used by Anvil (up to 3 per CPU, 10.4 GT/s) compared to Blacklight’s QuickPath Interconnect (QPI) links (up to 4 per CPU but at 6.4 GT/s). The newer interconnect links reduce inter-socket latency. Overall, the hierarchical load balancing model offers improvement in terms of memory access and runtime on larger machines, but this improvement is minimized on a newer, single shared memory node given its advancements in architecture compared to the older architecture utilized in [foteinos2014high].
7 Future Work
Figure 27 shows the latency of remote memory accesses and memory bandwidth on each of the supercomputers of our evaluation. We observe that latency increases as the size of a shared memory node increases (i.e., the number of cores). When transitioning from an Anvil shared memory node to Anvil’s distributed memory environment, average message latency between numa nodes increases from about 190 nanoseconds to 900 nanoseconds between distributed memory nodes while bandwidth significantly decreases (memory bandwidth within a single node vs. the network interconnect bandwidth between distributed nodes). These remote memory access latencies were obtained by running Intel’s version of the lmbench benchmark program [^4] https://github.com/intel/lmbench on a node within each supercomputer. Bandwidth specifications are also available for the CPUs utilized within Turing [TuringSpecs], Wahab [WahabSpecs], and Anvil [AnvilProcessorSpecs, AnvilSpecifications]. Given that our hierarchical load balancing model grants additional speedup even when executed on a shared memory node (i.e., Anvil) that has low latency and high bandwidth, the potential benefits of such a model are expected to be exacerbated where limiting the distance of data migration becomes paramount due to the much higher latency and lower network interconnect bandwidth in distributed memory. Although focused on generating structured hexahedral meshes for explosion shock wave simulations, such a model is implemented in [HierarchicalDistributed2025] using MPI-3 Remote Memory Access to create a virtual window for shared memory data transfer between distributed memory nodes. The method is able to utilize a static load balancing technique with low communication overhead because the neighborhood relationships of structured meshes remain fixed. Although overhead increases as the number of nodes utilized increases (due to implicit global synchronization caused by shared variables between nodes), the model indeed reduces average communication latency (by prioritizing local memory accesses) and improves the parallel efficiency of the method compared to when it utilizes traditional memory access techniques. This virtual window for shared memory access between distributed nodes could potentially be applied to anisotropic meshing; however, this is outside the scope of this paper and will be addressed in future work.
Figure 27: Shown is the remote memory access latency (collected using Intel’s lmbench program) and the bandwidth of each supercomputer [TuringSpecs, WahabSpecs, WahabSpecs2, AnvilProcessorSpecs, AnvilSpecifications].
Given its performance when adapting the aneurysm geometries at scaled complexities, we predict that CDT3D would be well suited for processing a geometry generated from an image of a large vascular structure (such as in Figure 5(a)). This would likely require many more elements compared to the smaller test cases of this study (e.g., millions to potentially hundreds of millions of elements which in the future may be needed for aneurysm cases requiring even finer mesh resolution such as microvascular aneurysms). This will be explored in future work. It should be noted that a routine within CDT3D, which attempts to remove boundary slivers [DrakopoulosCDT3D], was turned off for the scaled second aneurysm case. This likely contributed to the low quality elements seen in Figure 20(b). Both the original and optimized versions of the method sometimes crash when attempting to remove boundary slivers for this particular geometry at 50 million complexity. It should also be noted that although an Anvil shared memory node contains 128 cores, we tested CDT3D (and subsequently the other methods) with only 96 cores. CDT3D hangs during runtime when the method is executed with 128 cores. As these problems require further investigation, they will be remedied in future work. Nevertheless, the meshes generated by CDT3D for both large cases (and when utilizing up to 96 CPU cores) exhibit good overall quality. Additionally, slivers are indeed removed successfully when generating the smaller meshes (and when generating approximately 50 million elements for the carotid cavernous aneurysm case).